Given 3x+k-5 = kx-k+1, what is the value of k to make x=0 be the solution of the linear equation?
Answer by richwmiller(17219)

plug in 0 for x and find k
3x+k-5 = kx-k+1
k-5=-k+1
2k=6
k=3
check
plug in 3 for k and see if x=0
3x+3-5=3x-3+1
-2=-2
ok
when k=3 x can equal anything.